Solve:
a) t=5*logbase4 16
b) t=4*logbase1/3 3sqrt243

Answer by stanbon(75887)

Solve:
a) t=5*logbase4 16
t = 5*2
t = 10
--------------
b) t=4*logbase1/3 3sqrt243
t = 4* log base(1/3) 243^(1/3)
t = 4* log base(1/3) (3)^(5/3)
t = 4* log base(1/3) (1/3)^(-5/3)
t = 4*(-5/3)
t = -20/3
=============
Cheers,
Stan H.